How much money will have to put in the bank today to be able to withdraw $6,000 per year for the next 8 years assuming 12% intreat.

A 29,805
B 73,798
C 48,000
D 6,720
Answer #1

PMT =6000
Rate =12%
Number of Years =8
PV =PMT*((1-(1+r)^-n)/r =6000*((1-(1+12%)^-8)/12%) =29805

Option a is correct option